Say 40% of the class is female. What is the probability that 6 of the first 10 students walking in will be female?



1
Expert's answer
2022-05-23T23:34:36-0400

 Let "X=" the number of female: "X\sim Bin (n, p)."

Given "n=10, p=0.4, q=1-p=0.6."



"P(X=6)=\dbinom{10}{6}(0.4)^6(0.6)^{10-6}=0.111476736"